OpenStudy (anonymous):

g(x)=(-2)/(4m-2) f(x)=(2)/(m-4) Find: g(x)/f(x) HELP!!!!

OpenStudy (anonymous):

the variable here is x?

OpenStudy (anonymous):

or m?

myininaya (myininaya):

\[\frac{g(m)}{f(m)}=\frac{\frac{-2}{4m-2}}{\frac{2}{m-4}} =\frac{\frac{-2}{2(2m-1)}}{\frac{2}{m-4}}= \frac{\frac{-1}{2m-1}}{\frac{2}{m-4}} =\frac{-1}{2m-1} \cdot \frac{m-4}{2}\]